4. Solve the system of differential equations using elimination/substitution: x - 9y = 1 x+y = 4
Answer #1

Given question solved.Solve the system of diffrential equation using eliminatian/ substitution x - y = 1 x + y = 4 x-gy= x + y = 4 @ fon laminationNow Ausilony equation mt9=0 m-9 m = -6 m=231 Now, complementary sinal soluation Ye (t) = e Cosatte Pantiewer soluation is wheNow, substitute the value (9) value of y(t) into X+ D[ 9 Cosst + esinst - Yo] =4 9 Cossé + simok - Vy] =4 =>xt d at => x + [-